Solve

an+2 = 4an+1 + 21an; n  0 and a0 = 3; a1 = 4


1
Expert's answer
2021-02-25T01:51:36-0500

Given recurrence relation is-

"a_{n+2}=4a_{n+1}+21a_n" Where, "a_o=3,a_1=4"


The characterstics equation of the above relation is given by-


"x^2=4x+21\\\Rightarrow x^2-4x-21=0\\
\Rightarrow (x-7)(x+3)=0"


Then the roots of equation are "7,-3"

So ,

"a_n=c_1(7)^n+c_2(-3)^n~~~~~-(1)"


At "n=0"

"a_o=c_1+c_1\Rightarrow c_1+c_2=3~~~~~-(2)"


At "n=1"

"a_1=7c_1-3c_2\Rightarrow 7c_1-3c_2=4~~~-(3)"


On solving equation 2 and 3, we get-

"c_1=1.3,c_2=1.7"


Puuting the above values in eqs.(1)-

Hence eqs.(1)"\implies"


"a_n=(1.3)7^n+(1.7)(-3)^n"